# Evaluating small vision-language models as AI assistants for radio astronomical source analysis tasks

### v2

@@ 3.2 Model fine-tuning
The radio-llava model was trained using instruction fine-tuning on the Q&A radio dataset alone, as well as on the combined Q&A and caption datasets, starting from the pre-trained LLaVA-OneVision 7B model555lmms-lab/llava-onevision-qwen2-7b-ov. Keeping the vision encoder (siglip-so400m-patch14-384) frozen, we fine-tuned both the LLM (qwen2) and adapter (mlp2x_gelu) components using either full fine-tuning or the Low-Rank Adaptation (LoRA) (Hu et al, 2021) fine-tuning strategy [^6] LoRA is a lightweight training method that updates only small, low-rank matrices within the model instead of fine-tuning the entire model. This significantly reduces computational overhead and storage requirements while maintaining high performance. We set the LoRA rank and alpha scaling parameters to 64 and 128, respectively. The model was trained for either 1 epoch (shallow fine-tuning) or 3 epochs (deep fine-tuning). In all training runs, we used default hyperparameters, with a batch size of 1, a gradient accumulation step of 2, and a learning rate of 10-5.
@@ 3.3 Implementation details
Inference scripts developed for LLaVA and other VLM models tested in this study are provided in the following repository: https://github.com/SKA-INAF/radio-llava. This repository also includes a Streamlit [^9] https://streamlit.io/ application (see Fig. C.1 in the Appendix), allowing users to load a LLaVA model, upload an image, and interact with the assistant via a web interface. Table C in the Appendix provides an example of user-assistant conversations for two sample radio images, comparing responses from the base and fine-tuned LLaVA-OneVision 7B models.
@@ 4.1.1 Radio benchmarks
We used the radioimg-multilabel test dataset (5,718 images) to evaluate the models’ ability to detect extended or diffuse radio sources in input images. For this task, we applied the following prompt:
@@
Answer the question using the provided context (and examples). Report the identified class labels separated by commas, without any additional explanation text. Report just NONE if you cannot recognize any of the above classes in the image.
@@
We considered the rgz-smorph test dataset, containing $\sim$3,835 images from the VLA FIRST survey, each zoomed and centered around radio sources belonging to six distinct morphological classes: 1C-1P, 1C-2P, 1C-3P, 2C-2P, 2C-3P, 3C-3P. The model was evaluated using a single-label multi-class classification task with the following prompt:
@@
We used the radioimg-multilabel test dataset (5,718 images) to assess the models’ ability to identify radio sources with morphologies characteristic of extended radio galaxies. For this task, we applied the following prompt:
@@
Figure 2: Classification F1-scores obtained with VLMs of different sizes (0.5B, 2B, 3.1B, 7B, 8B, 72B) in zero-shot mode over B1-B6 evaluation benchmarks. We report the F1-score for individual classes, as well as the class-averaged F1-score (labelled as ’AVG’). LLaVA, TinyLLaVA, Qwen2VL, and InternVL models are respectively shown with blue, green, red, and orange histograms.
@@
We used the radioimg-multilabel test dataset (5,718 images) to evaluate the models’ ability to detect imaging artefacts in input images. For this task, we considered the following prompt:
@@
We used the Mirabest (Porter & Scaife, 2023) confident sample dataset, which contains 833 images from the VLA FIRST survey, each zoomed and centered around radio galaxies belonging to two distinct morphological classes: FR-I (397 images) and FR-II (436 images).
@@
FR-I: radio-loud galaxies characterized by a jet-dominated structure where the radio emissions are strongest close to the galaxys center and diminish with distance from the core;
FR-II: radio-loud galaxies characterized by a edge-brightened radio structure, where the radio emissions are more prominent in lobes located far from the galaxys core, with hotspots at the ends of powerful, well-collimated jets.
@@ 4.2 Zero-shot performance
In Fig. 2, we present the benchmark evaluation results for the base LLaVA-OneVision models (0.5B, 7B, 72B), shown in blue histograms, compared against alternative VLM models: TinyLLaVA 3.1B (green histogram), Qwen2VL models (2B, 7B, 72B) (red histograms), and InternVL models (2B, 8B) (orange histograms). For each benchmark, we reported the classification F1-score for individual classes, as well as the average F1-score across all classes (labeled as ’AVG’ in the plots).
@@
Figure 3: Classification F1-scores obtained with the radio-llava model on B1–B6 radio benchmarks, comparing fine-tuning on the Q&A training dataset (blue histograms) and the combined Q&A and caption datasets (orange histograms). For each training set, results are reported for different training strategies (full vs. LoRA fine-tuning) and training depths (shallow vs. deep). Results from the base model are shown as filled red histograms. The class-averaged F1-scores are labelled as ’AVG’.
@@ 4.3.1 Radio benchmarks
In Fig. 3 we report the classification F1-score of radio-llava fine-tuned models obtained on radio benchmarks for each class and overall (labelled as ’AVG’), compared to the base LLaVA-OneVision 7B model (solid red histograms). Blue histograms represent models fine-tuned on the Q&A dataset, using either deep/shallow full fine-tuning or LoRA fine-tuning. Orange histograms correspond to models fine-tuned on the combined Q&A and caption datasets.
@@
Overall, the achieved metrics remain well below those obtained using a pure vision encoder model specialized for each task. For instance, fine-tuning the SigLIP LLaVA encoder model (siglip-so400m-patch14-384) alone on the same multi-label, multi-class B1 classification task, using the same training and test data, resulted in an average F1-score of $\sim$80%. Notably, we achieved $\sim$60F1-score on diffuse sources, which multimodal models particularly struggle to detect, partly due to their underrepresentation in the training set (534 images). Similarly, for B2, previous work (Riggi et al., 2024c) trained a LightGBM classifier on features extracted solely from the SigLIP vision encoder, achieving an average F1-score of $\sim$74%.
@@
For B6, previous studies (Cecconello et al., 2024; Slijepcevic et al., 2024) have achieved over 90% classification accuracy using smaller fine-tuned encoders (ResNet18/ResNet50). However, those models were trained on an independent subset of the same survey data (MiraBest dataset, VLA survey) used for testing. In contrast, our work trained on different survey data with varying source/image size ratios—specifically, using ASKAP EMU pilot data for training and VLA zoomed-in source images for testing. This difference likely contributes to the poor results observed on B6.
@@ 4.3.2 Standard benchmarks
For comparison, Fig. 4 reports performance on non-radio benchmarks, using the same color labeling scheme of Fig. 3. Consistent with previous studies (Pan et al., 2024) specializing LLMs for astronomy, we observed a notable decline in model performance on previously learned tasks compared to the base model (solid red histograms). This task forgetting effect is particularly pronounced in full fine-tuning, becoming significant (more than a 20% accuracy drop) in deeper training runs. In line with findings by Biderman et al. (2024), LoRA fine-tuned models achieve lower performance on radio benchmarks but are more robust against task forgetting.
Catastrophic forgetting remains a critical challenge when fine-tuning LLMs. Recent studies (Zhai et al., 2023; Zhang et al., 2024) have analyzed this effect in multimodal models and proposed various strategies to mitigate it. One promising approach, successfully explored for language models by de Haan et al. (2024), involves expanding and curating the instruction-tuning dataset, followed by merging fine-tuned models with base models, using customizable balancing weights. In this work, we nearly doubled the size of our initial Q&A dataset, enriching it with more diverse image captions extracted from a large collection of scientific papers. As shown in Fig. 4, incorporating caption data (orange histograms) helped recover approximately 10 accuracy points across all standard benchmarks. This confirms trends observed in previous studies and underscores the importance of further curating our training dataset to enable future improvements. We attempted to merge our fully fine-tuned models with the base model using the MergeKit tool [^10] https://github.com/arcee-ai/mergekit. However, this approach remains largely experimental, as the tool currently supports merging only the language model component. While we extended its functionality to allow the merging of both the LLM and adapter components, we were unable to evaluate the capabilities of the merged model due to incompatibilities between the original LLaVA framework and the MergeKit required transformer library.
4.3.3 Impact of alternative hyperparameters
To assess the impact of the default training configuration on model performance, we fine-tuned the model on the Q&A dataset with alternative choices of selected hyperparameters, resulting in various model variants, which are labeled and summarized in Table 1.
@@ 4.3.3 Impact of alternative hyperparameters
Figure C.2 in the Appendix compares average metrics across all radio tasks for the original and variant models: the black solid histogram represents the original model, the black dashed histogram represents the original model with the alternative prompt, and the colored histograms correspond to the model variants. From the results, we conclude that the alternative configurations tested do not lead to performance improvements. Thus, the suboptimal performance is unlikely due to non-optimal hyperparameter selection but rather to dataset quality limitations. Furthermore, the reported metrics show minimal variation with the adopted prompt.
@@ 5 Summary
In this work, we investigated the feasibility of using small-scale Vision-Language Models (VLMs) as AI assistants for analyzing radio images, enabling tasks such as source classification, identification of specific object classes, and data exploration for quality assessment. Unlike conventional deep learning approaches, VLMs offer a more flexible, natural-language-driven interaction, reducing the need for complex coding or task-specific model adaptation. To this end, we fine-tuned LLaVA, a state-of-the-art VLM, on a custom dataset of over 59,000 radio images paired with instruction-based queries, along with an additional 38,000 image-caption pairs extracted from a large corpus of radio astronomical papers. The fine-tuning process leveraged both Q&A interactions and descriptive captions, enabling the model to handle a variety of radio analysis tasks, including source morphology classification, extended source detection, and artifact identification. The resulting radio-llava model was evaluated across six radio-specific benchmarks (B1--B6) and compared against baseline VLMs on non-astronomical multimodal tasks. Fine-tuned models and the developed software have been publicly released.
Our key findings can be summarized as follows:
@@
These findings highlight the potential of compact multimodal models for radio astronomy while also revealing key limitations that require further research to fully match the performance of specialized vision models. Future efforts should focus on improving vision-language alignment, curating larger, high-quality training datasets, and exploring hybrid fine-tuning strategies also for larger models ($\sim$70B) to mitigate task forgetting while maximizing domain-specific performance. Additionally, we plan to leverage the multi-image processing capabilities of the LLaVA-OneVision model for in-context learning of analyzed tasks. Future investigations will also explore its performance on new tasks requiring comparative analysis across multiple images, such as image retrieval of known source classes.
@@ A.1 Coarse-grained radio datasets
  • SARAO MeerKAT Galactic Plane Survey (SMGPS) (Goedhart et al., 2024): 2,704 images (14.2%)
  • ASKAP EMU main survey (Hopkins et al., 2024): 4,456 images (23.4%)
  • ASKAP EMU pilot survey (Norris et al., 2021): 5,860 images (30.7%)
  • ASKAP EMU pilot Galactic Plane surveys (Umana et al., 2021): 6,040 images (31.7%)
@@
More than one label can be assigned to each image, depending on the object/features the user recognizes in the image. The number of images that have been assigned each specific label is reported in Table 2.
@@
Table 2: The number of images in the radioimg-multilabel dataset that have been assigned each specific label. Multiple labels can be assigned to a single image, as they are not mutually exclusive.
@@
The dataset currently includes a collection of 9,570 radio images extracted from the VLA FIRST survey (Becker et al., 1995) and annotated in the Radio Galaxy Zoo (RGZ) crowdsourced project (Banfield et al., 2015). Each image is centered and zoomed on radio sources of 6 different morphological classes, defined on the basis of the observed number of components (C) and peaks (P) as follows: 1C-1P, 1C-2P, 1C-3P, 2C-2P, 2C-3P, 3C-3P. The entire dataset was split into two samples. The first one, containing 5,735 images ($\sim$1000 per class), was used to create the conversational dataset, while the remaining sample (3,835 images, $\sim$600 per class) was reserved for model evaluation scopes.
@@
The dataset currently includes a collection of 17,062 radio images extracted from the SMGPS survey (Goedhart et al., 2024), each centered and zoomed [^12] The original image crop size is set to 1.5 times the size of the source bounding box. on radio sources listed in the SMGPS extended source catalogue (Bordiu et al., 2025). This includes single- or multi-island sources with morphologies classified as: EXTENDED or DIFFUSE. Furthermore, a fraction of the catalogued sources also have an astronomical class label obtained either through morphological considerations or cross-matching with various Galactic source catalogues (see Bordiu et al. 2025 for details). Available class labels are: GALAXY (radio galaxy), HII (Hii region), PN (planetary nebula), SNR (supernova remnant), PULSAR (pulsar), STAR (generic radio star), YSO (young stellar objects), LBV (luminous blue variable star), WR (Wolf-Rayet star), HMXB (high-mass X-ray binary), LMXB (low-mass X-ray binary). Sources cross-matching to multiple catalogues have more than one label assigned. All the above source annotations are taken into account to generate the conversational dataset.

### v3

@@ 3.2 Model fine-tuning
The radio-llava model was trained using instruction fine-tuning on the Q&A radio dataset alone, as well as on the combined Q&A and caption datasets, starting from the pre-trained LLaVA-OneVision 7B model555lmms-lab/llava-onevision-qwen2-7b-ov. Keeping the vision encoder (siglip-so400m-patch14-384) frozen, we fine-tuned both the LLM (qwen2) and adapter (mlp2x_gelu) components using either full fine-tuning or the Low-Rank Adaptation (LoRA) Hu et al (2021) fine-tuning strategy [^6] LoRA is a lightweight training method that updates only small, low-rank matrices within the model instead of fine-tuning the entire model. This significantly reduces computational overhead and storage requirements while maintaining high performance. We set the LoRA rank and alpha scaling parameters to 64 and 128, respectively. The model was trained for either 1 epoch (shallow fine-tuning) or 3 epochs (deep fine-tuning). In all training runs, we used default hyperparameters, with a batch size of 1, a gradient accumulation step of 2, and a learning rate of 10-5.
@@ 3.3 Implementation details
Inference scripts developed for LLaVA and other VLM models tested in this study are provided in the following repository: https://github.com/SKA-INAF/radio-llava. This repository also includes a Streamlit [^9] https://streamlit.io/ application (see Figure C.1 in the Appendix), allowing users to load a LLaVA model, upload an image, and interact with the assistant via a web interface. Table C in the Appendix provides an example of user-assistant conversations for two sample radio images, comparing responses from the base and fine-tuned LLaVA-OneVision 7B models.
@@ 4.1.1 Radio benchmarks
We used the radioimg-multilabel test dataset (5,718 images) to evaluate the models’ ability to detect extended or diffuse radio sources in input images.
For this task, we applied the following prompt:
@@
Answer the question using the provided context (and examples). Report the identified class labels separated by commas, without any additional explanation text. Report just NONE if you cannot recognise any of the above classes in the image.
@@
We considered the rgz-smorph test dataset, containing $\sim$3,835 images from the VLA FIRST survey, each zoomed and centred around radio sources belonging to six distinct morphological classes: 1C-1P, 1C-2P, 1C-3P, 2C-2P, 2C-3P, 3C-3P.
The model was evaluated using a single-label multi-class classification task with the following prompt:
@@
We used the radioimg-multilabel test dataset (5,718 images) to assess the models’ ability to identify radio sources with morphologies characteristic of extended radio galaxies.
For this task, we applied the following prompt:
@@
Figure 2: Classification F1-scores obtained with VLMs of different sizes (0.5B, 2B, 3.1B, 7B, 8B, 72B) in zero-shot mode over B1-B6 evaluation benchmarks. We report the F1-score for individual classes, as well as the class-averaged F1-score (labelled as ’AVG’). LLaVA, TinyLLaVA, Qwen2VL, and InternVL models are respectively shown with blue, green, red, and orange histograms. OpenAI GPT4.1 model is shown with black histograms.
@@
We used the radioimg-multilabel test dataset (5,718 images) to evaluate the models’ ability to detect imaging artefacts in input images.
For this task, we considered the following prompt:
@@
We used the Mirabest Porter & Scaife (2023) confident sample dataset, which contains 833 images from the VLA FIRST survey, each zoomed and centred around radio galaxies belonging to two distinct morphological classes: FR-I (397 images) and FR-II (436 images).
@@
FR-I: radio-loud galaxies characterised by a jet-dominated structure where the radio emissions are strongest close to the galaxys centre and diminish with distance from the core;
FR-II: radio-loud galaxies characterised by an edge-brightened radio structure, where the radio emissions are more prominent in lobes located far from the galaxys core, with hotspots at the ends of powerful, well-collimated jets.
@@ 4.2 Zero-shot performance
We evaluated the zero-shot performance of LLaVA models of varying sizes on radio benchmarks, comparing with alternative open-weight VLMs and a representative commercial model (OpenAI GPT 4.1). Results are reported in Figure 2 and discussed in the following paragraphs. For each benchmark, we report the classification F1-score for individual classes, as well as the average F1-score across all classes (labelled as ’AVG’ in the plots).
4.2.1 Open-weight models
In Figure 2, we present the benchmark evaluation results for the base LLaVA-OneVision models (0.5B, 7B, 72B), shown in blue histograms, compared against alternative open-weight VLM models: TinyLLaVA 3.1B (green histogram), Qwen2VL models (2B, 7B, 72B) (red histograms), and InternVL models (2B, 8B) (orange histograms).
@@ 4.2.1 Open-weight models
Figure 3: Classification F1-scores obtained with the radio-llava model on B1–B6 radio benchmarks, comparing fine-tuning on the Q&A training dataset (blue histograms) and the combined Q&A and caption datasets (orange histograms). For each training set, results are reported for different training strategies (full vs. LoRA fine-tuning) and training depths (shallow vs. deep). Results from the base model are shown as filled red histograms. Results obtained with a fine-tuned vision-only model (siglip-so400m-patch14-384 encoder) are shown as filled black histograms. The class-averaged F1-scores are labelled as ’AVG’.
4.2.2 Commercial closed-weight models
Performing a comprehensive end-to-end benchmark evaluation across major proprietary solutions (e.g. OpenAI GPT, Google Gemini, Anthropic Claude) is not straightforward, as it would require academic institution to enter into contractual agreements with private providers to cover the cost of executing benchmarks via their APIs. Unlike public user interfaces, these APIs typically operate under separate pricing and access tiers. Nonetheless, we recognise the value of such an analysis for understanding the feasibility and cost-effectiveness of commercial LLM APIs in scientific benchmarking. Therefore, we made an effort to evaluate at least one commercial model - GPT-4.1 via the OpenAI API platform.
Benchmarks were split into multiple sub-tasks of approximately 500 images each to stay within the maximum batch file size limit (200 MB), with each mini-batch consisting of 80K--230K input tokens, costing around 0.23\$ - implying a total of $\sim$2.5\$ per benchmark and under 20\$ for the full suite. As Tier 1 users, we were able to run one or two mini-batches per day without exceeding the token-per-day (TPD) limit of 900,000 tokens.
The GPT-4.1 benchmark results, shown in Figure 2 as black histograms, indicate superior performance in tasks B1-B3 and B6. In tasks B1 and B6, GPT-4.1 outperforms all open-weight models by a substantial margin -- approximately 20% in average classification score. For tasks B2 and B3, the improvement is more modest, generally below 10%. Interestingly, GPT-4.1 underperforms in tasks B4 and B5, where its classification metrics fall below those of several open-weight models. These results may reflect both the advantage conferred by GPT-4’s significantly larger parameter count [^10] Notably, GPT-4 models are estimated to have approximately 25 times more parameters (around 1.8 trillion parameters from various sources) than the largest open-weight models evaluated in this work and broader pretraining corpus, as well as limitations in its exposure to domain-specific astronomical concepts or visual patterns. While GPT-4.1 currently achieves the best overall performance in our benchmarks, the relatively small gap in several tasks -- combined with the flexibility, transparency, and lower deployment costs of open-weight models -- suggests there remains meaningful room for their development and application in specialized astronomical workflows.
@@ 4.3.1 Radio benchmarks
In Figure 3 we report the classification F1-score of radio-llava fine-tuned models obtained on radio benchmarks for each class and overall (labelled as ’AVG’), compared to the base LLaVA-OneVision 7B model (solid red histograms). Blue histograms represent models fine-tuned on the Q&A dataset, using either deep/shallow full fine-tuning or LoRA fine-tuning. Orange histograms correspond to models fine-tuned on the combined Q&A and caption datasets. For comparison, we also report baseline metrics (shown as black histograms) obtained using a vision-only classifier that shares the same vision encoder as the LLaVA model (siglip-so400m-patch14-384). This classifier was fine-tuned and evaluated on the same training and test datasets.
@@
Overall, the achieved metrics remain well below those obtained using a vision-only model specialized for each task, which consistently reaches over 80--85% accuracy across all benchmarks - even after just 10 training epochs. The performance gap is especially notable in task B1, where the vision-only model attains an F1-score of approximately 60% for diffuse sources -- class that multimodal models tend to struggle with, likely due to their underrepresentation in the training set (only 534 images). Similarly, in task B2, the model achieves over 80% accuracy across all morphological classes. These results also surpass our previous baseline of 74% average F1-score Riggi et al. (2024c), which was obtained by training a LightGBM classifier on features extracted solely from the SigLIP vision encoder. At present, a fair comparison between the vision-only and LLaVA models cannot be provided for benchmark B6 due to the lack of a shared training dataset. Specifically, the FR-I/FR-II labels used in the LLaVA training were derived from object detection conversations on wide-field images, whereas training a vision-only classifier would require centred cutouts around FR-I/FR-II sources. Previous FR-I/FR-II classification studies Cecconello et al. (2024); Slijepcevic et al. (2024) have achieved over 90% classification accuracy using smaller fine-tuned encoders (ResNet18/ResNet50). However, those models were trained on an independent subset of the same survey data (MiraBest dataset, VLA survey) used for testing. In contrast, our work trained on different survey data with varying source/image size ratios -- specifically, using ASKAP EMU pilot data for training and VLA zoomed-in source images for testing. This difference likely contributes to the poor results observed on B6. We plan to update the dataset accordingly in future work to enable a consistent evaluation across both models.
@@ 4.3.2 Standard benchmarks
For comparison, Figure 4 reports performance on non-radio benchmarks, using the same colour labelling scheme of Figure 3. Consistent with previous studies Pan et al. (2024) specializing LLMs for astronomy, we observed a notable decline in model performance on previously learned tasks compared to the base model (solid red histograms). This task forgetting effect is particularly pronounced in full fine-tuning, becoming significant (more than a 20% accuracy drop) in deeper training runs. In line with findings by Biderman et al. (2024), LoRA fine-tuned models achieve lower performance on radio benchmarks but are more robust against task forgetting.
Catastrophic forgetting remains a critical challenge when fine-tuning LLMs. Recent studies Zhai et al. (2023); Zhang et al. (2024) have analysed this effect in multimodal models and proposed various strategies to mitigate it. One promising approach, successfully explored for language models by de Haan et al. (2024), involves expanding and curating the instruction-tuning dataset, followed by merging fine-tuned models with base models, using customizable balancing weights [^11] The merging tool used in de Haan et al. (2024) -- MergeKit https://github.com/arcee-ai/mergekit -- currently supports merging only the LLM components and requires extension to also include the LLaVA adapter and vision components.. In this work, we nearly doubled the size of our initial Q&A dataset, enriching it with more diverse image captions extracted from a large collection of scientific papers. As shown in Figure 4, incorporating caption data (orange histograms) helped recover approximately 10 accuracy points across all standard benchmarks. This confirms trends observed in previous studies and underscores the importance of further curating our training dataset to enable future improvements.
4.3.3 Diagnostic analysis
To assess the impact of the default training configuration on model performance, we fine-tuned the model on the Q&A dataset with alternative choices of selected hyperparameters, resulting in various model variants, which are labelled and summarised in Table 1.
@@ 4.3.3 Diagnostic analysis
Figure C.2 in the Appendix compares average metrics across all radio tasks for the original and variant models: the black solid histogram represents the original model, the black dashed histogram represents the original model with the alternative prompt, and the coloured histograms correspond to the model variants. From the results, we conclude that the alternative configurations tested do not lead to performance improvements. Thus, the suboptimal performance is unlikely due to non-optimal hyperparameter selection but rather to dataset quality limitations. Furthermore, the reported metrics show minimal variation with the adopted prompt.
To assess the impact of using automated data curation via InternVL, we initially conducted experiments using a fully templated, uncurated dataset without InternVL-generated variation. While benchmark performance metrics remained broadly similar, we observed a marked degradation in the model outputs, with responses frequently mirroring the rigid structure of the templates. This motivated us to setup an automated data curation to enhance linguistic diversity and model conversational ability.
@@ 5 Summary
In this work, we investigated the feasibility of using small-scale Vision-Language Models (VLMs) as AI assistants for analysing radio images, enabling tasks such as source classification, identification of specific object classes, and data exploration for quality assessment. Unlike conventional deep learning approaches, VLMs offer a more flexible, natural-language-driven interaction, reducing the need for complex coding or task-specific model adaptation. To this end, we fine-tuned LLaVA, a state-of-the-art VLM, on a custom dataset of over 59,000 radio images paired with instruction-based queries, along with an additional 38,000 image-caption pairs extracted from a large corpus of radio astronomical papers. The fine-tuning process leveraged both Q&A interactions and descriptive captions, enabling the model to handle a variety of radio analysis tasks, including source morphology classification, extended source detection, and artifact identification. The resulting radio-llava model was evaluated across six radio-specific benchmarks (B1--B6) and compared against baseline VLMs on non-astronomical multimodal tasks. Fine-tuned models and the developed software have been publicly released.
Our key findings can be summarised as follows:
@@
These findings highlight the potential of compact multimodal models for radio astronomy while also revealing key limitations that require further research to fully match the performance of specialized vision models. Future efforts should focus on improving vision-language alignment, curating larger, high-quality training datasets, and exploring hybrid fine-tuning strategies also for larger models ($\sim$70B) to mitigate task forgetting while maximizing domain-specific performance. Additionally, we plan to leverage the multi-image processing capabilities of the LLaVA-OneVision model for in-context learning of analysed tasks. Future investigations will also explore its performance on new tasks requiring comparative analysis across multiple images, such as image retrieval of known source classes.
@@ A.1 Coarse-grained radio datasets
  • SARAO MeerKAT Galactic Plane Survey (SMGPS) Goedhart et al. (2024): 2,704 images (14.2%)
  • ASKAP EMU main survey Hopkins et al. (2025): 4,456 images (23.4%)
  • ASKAP EMU pilot survey Norris et al. (2021): 5,860 images (30.7%)
  • ASKAP EMU pilot Galactic Plane surveys Umana et al. (2021): 6,040 images (31.7%)
@@
More than one label can be assigned to each image, depending on the object/features the user recognises in the image. The number of images that have been assigned each specific label is reported in Table A1.
@@
Table A1: The number of images in the radioimg-multilabel dataset that have been assigned each specific label. Multiple labels can be assigned to a single image, as they are not mutually exclusive.
@@
The dataset currently includes a collection of 9,570 radio images extracted from the VLA FIRST survey Becker et al. (1995) and annotated in the Radio Galaxy Zoo (RGZ) crowdsourced project Banfield et al. (2015). Each image is centred and zoomed on radio sources of 6 different morphological classes, defined on the basis of the observed number of components (C) and peaks (P) as follows: 1C-1P, 1C-2P, 1C-3P, 2C-2P, 2C-3P, 3C-3P. The entire dataset was split into two samples. The first one, containing 5,735 images ($\sim$1000 per class), was used to create the conversational dataset, while the remaining sample (3,835 images, $\sim$600 per class) was reserved for model evaluation scopes.
@@
The dataset currently includes a collection of 17,062 radio images extracted from the SMGPS survey Goedhart et al. (2024), each centred and zoomed [^13] The original image crop size is set to 1.5 times the size of the source bounding box. on radio sources listed in the SMGPS extended source catalogue Bordiu et al. (2025). This includes single- or multi-island sources with morphologies classified as: EXTENDED or DIFFUSE. Furthermore, a fraction of the catalogued sources also have an astronomical class label obtained either through morphological considerations or cross-matching with various Galactic source catalogues (see Bordiu et al. (2025) for details). Available class labels are: GALAXY (radio galaxy), HII (Hii region), PN (planetary nebula), SNR (supernova remnant), PULSAR (pulsar), STAR (generic radio star), YSO (young stellar objects), LBV (luminous blue variable star), WR (Wolf-Rayet star), HMXB (high-mass X-ray binary), LMXB (low-mass X-ray binary). Sources cross-matching to multiple catalogues have more than one label assigned. All the above source annotations are taken into account to generate the conversational dataset.